OpenStudy (anonymous):

A tricycle has a tire with a circumference of 22 inches. How many revolutions does the tire make as the tricycle rolls 11 feet?

OpenStudy (anonymous):

Distance / Circumference = (11*12) / 22 = 132 / 22 = 6 revolutions

OpenStudy (anonymous):

Where did you get 12?

OpenStudy (anonymous):

12 inches per foot

OpenStudy (anonymous):

Got it. Thanks.